# Natural Language Actor-Critic: Scalable Off-Policy Learning in Language Space

### v1

@@ 8 Discussion
In this paper, we propose NLAC, a new actor-critic algorithm for training LLM agents under the paradigm of learning values in natural language space. In our work, the natural language values not only comment on the optimality of an action, but also articulate why by predicting and analyzing future outcomes. The key innovation we propose to enable this is a novel language Bellman backup that trains language successor function to generate possible future rollouts using only one-step samples obtained off-policy. Then, an LLM policy can be improved by refining its own suboptimal actions. This procedure improves one of the main challenges of RL fine-tuning for complex tasks—reliance on random exploration to uncover better actions–and significantly improves sample efficiency. Empirically, we show that NLAC greatly outperforms other prompting and fine-tuning baselines on long-horizon tasks involving dialogue and tool-use. As future work, we aim to see if our approach can be combined with traditional RL policy optimization by extracting a generative scalar value from our critiques to enable more sophisticated policy improvement strategies.
Appendix A Proofs
A.1 Proof of Theorem 5.1
@@ Appendix A Proofs
Since $\mathcal{T}_{\Phi}^{\pi}$ is contraction, the Banach Fixed-Point Theorem with Assumption 5.2 guarantees that minimization of our training objective in Equation 2 results in the unique fixed point
@@
Next, we want to establish a monotonic mapping from representations $\Phi^{\pi}(s,a)$ to true Q-values $Q^{\pi}(s,a)$. The true $Q$-function is defined as $Q^{\pi}(s_{t},a_{t})=\mathbb{E}_{\pi}[\sum_{\tau=t}^{\infty}\gamma^{\tau-t}r(s_{\tau},a_{\tau})]$. Using Assumption 5.1, and substituting the fixed-point result from earlier, we have:
@@
A.2 Proof of Theorem 5.2
First, we show monotonic policy improvement in each iteration of our algorithm. Recall that the refinement policy $\pi^{r}$ is defined as the oracle that generates $a^{r}$ such that $Q_{L}^{\pi}(s,a^{r})\geq Q_{L}^{\pi}(s,a)$. By Theorem 5.1, a monotonic mapping $g$ exists between $Q_{L}^{\pi}$ and $Q^{\pi}$, and ensures that the ordering established in the language space holds in the true scalar space:
@@
Note that the policy improvement objective in Equation 4 updates $\pi$ to $\pi^{\prime}$, which is an approximation of the refinement policy $\pi^{r}$. By definition of refinement policy, this guarantees that the new policy $\pi^{\prime}$ is monotonically better than $\pi$ in terms of expected return: $Q^{\pi^{\prime}}(s,a)\geq Q^{\pi}(s,a)\ \text{for all }s,a$.
@@ B.4 Training Details
Our fine-tuning baselines were implemented using the Volcano Engine Reinforcement Learning (verl) library (sheng2024hybridflow). We train on $8$ H20 GPU nodes, resulting in $64$ GPUs total, for a total of $30,720$ gradient steps. Training took <48<48 hours for each benchmark. We used the following hyperparameter configuration for each benchmark, after some minimal amount of tuning:
@@ B.5 Training the Refinement Policy
To train the refinement policy, we adopt a similar approach as kumar2024traininglanguagemodelsselfcorrect did to train policies to self-correct, but generalized to multi-step MDPs. Namely, for some on-policy sample $s_{t},a_{t},a_{t}^{r}$ from the refinement policy, we additionally train on the loss:
@@ B.6 Mitigating Catastrophic Forgetting
In our experiments, we choose to stop training before catastrophic forgetting became a noticeable issue. However, we also considered methods to minimize forgetting, notably by incorporating learning without forgetting into our training objective  (li2016learningwithoutforgetting). The way we do so is by introducing auxiliary loss functions that penalizes divergence from the initial, pre-trained LLM. We introduce two additional losses for the language Bellman backup and refinement policy, respectively:

### v2

@@ 7 Discussion
In this paper, we propose NLAC, a new actor-critic algorithm for training LLM agents under the paradigm of learning values in natural language space. In our work, the natural language values not only comment on the optimality of an action, but also articulate why by predicting and analyzing future outcomes. The key innovation we propose to enable this is a novel language Bellman backup that trains language successor function to generate summaries of future rollouts using only one-step samples obtained off-policy. Then, an LLM policy can be improved by refining its own suboptimal actions. Empirically, we show that NLAC greatly outperforms other prompting and fine-tuning baselines on long-horizon tasks involving dialogue and tool-use. One apparent limitation of our approach is the reliance on the base LLM’s ability to both generate plausible futures and refine its own actions through carefully tuned prompts, which can break when tasks become prohibitively complex. An important direction for future research is the removal of this dependency, potentially by adding the ability to decompose complex tasks hierarchically into manageable components.
Appendix A Theoretical Analysis
The goal of this section is to show that policy iteration using our proposed NLAC method converges to the optimal policy. For traditional actor-critic algorithms, this involves showing (1) convergence of learned Q-values via the Bellman backup, and (2) monotonic improvement of the trained policy. However, such analysis does not apply because the Q-values we learn are textual rather than scalar.
Assumption A.1.
Assumption A.2.
For any state $s$ and action $a$ , the language Bellman backup satisfies
Theorem A.3.
Consider policy evaluation via Equation 4.1 and let $Q_{L}^{\pi}$ be the natural language critic at convergence. For any state $s$ and action $a$, there exists monotonic mapping $g$ such that $Q^{\pi}(s,a)=g(Q_{L}^{\pi}(s,a))$, where $Q^{\pi}$ denotes the true scalar Q-function.
Our main result makes a precise connection between our learned critic and the true Q-function. The crux of our proof involves showing that our training objective results in a fixed point where $\Phi^{\pi}(s,a)$ are successor features of the underlying MDP (Barreto et al., 2017).
During policy improvement, we update the policy towards a refinement policy $\pi^{r}$. By definition, the refinement policy is an oracle that for any state $s$ and action $a$, generates $a^{r}$ such that $Q_{L}(s,a^{r})\geq Q_{L}(s,a)$. Combining this with the result of Theorem A.3, we arrive at our second main result:
Theorem A.4.
Repeated application of policy evaluation via Equation 4.1 and policy improvement via Equation 4.2 from a policy $\pi_{0}$ converges to policy $\pi^{*}$ such that $Q^{\pi^{*}}(s,a)\geq Q^{\pi}(s,a)$ for any state $s$ and action $a$, and other policy $\pi$.
Hence, we are able to show that under the aforementioned assumptions, our approach NLAC can provably find the optimal policy for an underlying MDP.
A.1 Proof of Theorem A.3
@@ Appendix A Theoretical Analysis
Since $\mathcal{T}_{\Phi}^{\pi}$ is contraction, the Banach Fixed-Point Theorem with Assumption A.2 guarantees that minimization of our training objective in Equation 4.1 results in the unique fixed point
@@
Next, we want to establish a monotonic mapping from representations $\Phi^{\pi}(s,a)$ to true Q-values $Q^{\pi}(s,a)$. The true $Q$-function is defined as $Q^{\pi}(s_{t},a_{t})=\mathbb{E}_{\pi}[\sum_{\tau=t}^{\infty}\gamma^{\tau-t}r(s_{\tau},a_{\tau})]$. Using Assumption A.1, and substituting the fixed-point result from earlier, we have:
@@
A.2 Proof of Theorem A.4
First, we show monotonic policy improvement in each iteration of our algorithm. Recall that the refinement policy $\pi^{r}$ is defined as the oracle that generates $a^{r}$ such that $Q_{L}^{\pi}(s,a^{r})\geq Q_{L}^{\pi}(s,a)$. By Theorem A.3, a monotonic mapping $g$ exists between $Q_{L}^{\pi}$ and $Q^{\pi}$, and ensures that the ordering established in the language space holds in the true scalar space:
@@
Note that the policy improvement objective in Equation 4.2 updates $\pi$ to $\pi^{\prime}$, which is an approximation of the refinement policy $\pi^{r}$. By definition of refinement policy, this guarantees that the new policy $\pi^{\prime}$ is monotonically better than $\pi$ in terms of expected return: $Q^{\pi^{\prime}}(s,a)\geq Q^{\pi}(s,a)\ \text{for all }s,a$.
@@ B.4 Training Details
Our fine-tuning baselines were implemented using the Volcano Engine Reinforcement Learning (verl) library (Sheng et al., 2024). We train on $8$ H20 GPU nodes, resulting in $64$ GPUs total, for a total of $30,720$ gradient steps. Training took <48<48 hours for each benchmark. We used the following hyperparameter configuration for each benchmark, after some minimal amount of tuning:
@@
In practice, we found it helpful to implement a prioritized replay buffer weighted by $\mathcal{L}_{1}(s_{t},a_{t},s_{t+1})$ with sampling parameter $\alpha$ (Schaul et al., 2016). This is because in many tasks, though a base LLM policy may achieve low reward in a large proportion of rollouts, many actions in these unsuccessful rollouts are still optimal. Therefore, to improve learning efficiency, we prioritize training on samples where the agent is likely to take a suboptimal action, using critic loss as a proxy for the likelihood.
@@ B.5 Training the Refinement Policy
To train the refinement policy, we adopt a similar approach as Kumar et al. (2024) did to train policies to self-correct, but generalized to multi-step MDPs. Namely, for some on-policy sample $s_{t},a_{t},a_{t}^{r}$ from the refinement policy, we additionally train on the loss:
@@ B.6 Mitigating Catastrophic Forgetting
In our experiments, we choose to stop training before catastrophic forgetting became a noticeable issue. However, we also considered methods to minimize forgetting, notably by incorporating learning without forgetting into our training objective  (Li & Hoiem, 2016). The way we do so is by introducing auxiliary loss functions that penalizes divergence from the initial, pre-trained LLM. We introduce two additional losses for the language Bellman backup and refinement policy, respectively: